OpenStudy (anonymous):

what is C so that y=2x+3 is tangent to the curve y+x^2 +C?

OpenStudy (raden):

i think that's y=x^2 +C, right ?

OpenStudy (raden):

if yes, first take y1=y2, we get a quadratic equation and then to find value of c, take D=0 (D=discriminant) solve for c?

OpenStudy (anonymous):

yes, thats what I was thinking, do they mean y=x^2 +C

OpenStudy (raden):

did u get it?

OpenStudy (anonymous):

y=2x+3...........i y=x^2+C........ii

OpenStudy (anonymous):

Now, substitute eqn i in eqn ii and tell me what u get

OpenStudy (anonymous):

If you want to see the calculation look at the attached file.

OpenStudy (anonymous):

thank you very much guys for the help! can I ask why is the D=0 (D=discriminant)?

OpenStudy (amorfide):

descriminant is 0 because it is a tangent to the curve, when the descriminant is 0 it has a repeated root, so it just touches the curve, if it were to be greater than 0 it would have two roots which means it would no longer be a tangent, if it were less than 0 it would have two complex roots, not being a tangent,